### v1

@@ Abstract
Prompt-tuning has demonstrated parameter-efficiency in fusing unimodal foundation models for multimodal tasks. However, its limited adaptivity and expressiveness lead to suboptimal performance when compared with other tuning methods. In this paper, we address this issue by disentangling the vanilla prompts to adaptively capture dataset-level and instance-level features. Building upon this disentanglement, we introduce the mixture of prompt experts (MoPE) technique to enhance expressiveness. MoPE leverages multimodal pairing priors to route the most effective prompt on a per-instance basis. Compared to vanilla prompting, our MoPE-based conditional prompting exhibits greater expressiveness for multimodal fusion, scaling better with the training data and the overall number of trainable parameters. We also study a regularization term for expert routing, leading to emergent expert specialization, where different experts focus on different concepts, enabling interpretable soft prompting. Extensive experiments across three multimodal datasets demonstrate that our method achieves state-of-the-art results, matching or even surpassing the performance of fine-tuning, while requiring only 0.8% of the trainable parameters. Code will be released: https://github.com/songrise/MoPE.
Keywords:
Multimodal Fusion Prompt Tuning Mixture of Experts
1 Introduction
@@
Figure 1: High-level motivation of MoPE-based conditional prompting. a) vanilla prompt tuning learn a global shared prompt for all instances, which may not be optimal for each instance; b) MoPE-based conditional prompt tuning optimize multiple specialized prompt experts and is instance-wise adaptative.
Empowered with billion-scale data and highly scalable model architectures, recent unimodal pre-trained models [35, 28, 3, 4, 17], also known as foundation models, have demonstrated their transferable power to various downstream tasks [26, 30, 51]. However, transferring multimodal foundation models for downstream applications is less flexible. Recent explorations, such as CLIP [34], employ two-tower designs that entangle the two encoders, meaning that replacing either one would necessitate expensive paired pre-training from scratch. This limitation hinders the broader application that would benefit from a specific unimodal architecture, such as longer context windows. To democratize foundation models, a compelling question arises: How can we efficiently combine separately pretrained unimodal foundation models for downstream multimodal tasks?
Recently, prompt tuning [20, 11] has emerged as a novel paradigm for parameter-efficient transfer learning. While originally proposed for low-shot adaptation to tasks with the same modality as they were pretrained on [20, 11, 50], it has also proven effective for multimodal fusion. Typically, this is achieved by abstracting the representation from one modality into one or a few prompts, which are fed into the frozen Transformer of another modality for fusion [39, 22, 21, 45].
@@
Despite the effectiveness of prompt tuning in inter-modality transfer learning, directly adopting it for multimodal fusion could yield suboptimal results, particularly due to its limited adaptivity and scalability. Notably, while prompt tuning generally performs well in low-data regimes, it can be less effective when applied to full-shot training on the entire dataset with a challenging objective [7, 22, 39, 44, 19]. This reduced efficacy could be attributed to at least two factors. First, the vanilla prompt tuning used in previous prompt-based fusion methods [39, 21, 22] is not designed for multimodal tasks. These methods simply optimize a globally shared prompt for all instances [11, 20, 39], neglecting the multimodal interplay and potentially failing to capture the locally optimal representation for each input instance. Second, the relatively small number of trainable parameters (compared with fine-tuning) can lead to underfitting in multimodal datasets with a long-tail distribution and complex cross-modal mapping [27, 32, 42].
@@
To address these challenges, increasing the number of learnable prompts, known as “length-scaling,” appears to be a plausible solution. Despite its apparent simplicity, a lot of research indicates that the performance gains from length-scaling quickly reach saturation in both unimodal [7, 19] and multimodal cases [22, 39, 44]. Furthermore, over-length prompts may even lead to worse results [11, 44, 14, 20, 16]. Recent theoretical analyses have substantiated these empirical observations, highlighting the limited expressiveness of prompt tuning [42, 32]. A direct implication of their analysis is that simply increasing prompt length is insufficient to address the issues of adaptivity and expressiveness.
Based on the above observations, we aim to enhance the expressiveness of the vanilla prompt used in multimodal fusion to improve adaptivity and scalability. Our motivation is to utilize the multimodal input pairing information as a prior, and to scale up the prompt in another dimension—the number of prompt experts—rather than length, as illustrated in Fig. 1. To accomplish this, we first propose a conditional prompt tuning method, which disentangles the vanilla global-shared prompt vector into static and dynamic prompts to adaptively capture global and local features. The cornerstone of our dynamic prompt generation is our innovative MoPE (Mixture of Prompt Experts) design. In this design, we fix the prompt length used in attention and scale up expressiveness by learning multiple prompt experts and a router at each layer. We then use the representation from another modality as a prior to route the most effective dynamic prompt for each instance, enabling adaptive fusion.
We show that our MoPE-based fusion method alleviates the two issues in previous prompt-based fusion methods by allowing instance-wise adaptive prompting and avoiding over-length deterioration for better scalability. We systematically compare our MoPE-based fusion method with previous fusion methods on three multimodal datasets, UPMC Food-101 [41], SNLI-VE [43] and MM-IMDB [1]. Not only does our method deliver state-of-the-art results on all datasets, it also has higher parameter-efficiency compared with methods based on vanilla prompting, due to our improved adaptivity and expressiveness. Specifically, we achieve results comparable to fine-tuning while requiring only $\sim 0.8\%$ of the trainable parameters, and we outperform previous state-of-the-art (SOTA) prompt-based methods by a significant margin. Additionally, we demonstrate the superior scalability of our method. Our analysis indicates that increasing the number of experts (i.e. “expert-scaling”) tends to be a better alternative than length-scaling, which avoids performance deterioration with an overly-long prompt. In addition to parameter scaling, our method also demonstrates better scalability with increased training data. Finally, as the first endeavor to use prompts as experts, we analyze how different methods of expert routing can affect expert specialization. Thanks to the proposed regularization term, we observe the emergence of specialized prompt experts after end-to-end training, achieving high interpretability compared to black-box conditional prompting. We summarize our contributions as follows:
  • We design a conditional prompt tuning method for multimodal fusion, which disentangles the vanilla prompt as static, dynamic and mapped prompts for better adaptiveness.
@@
  • A combination of regularization terms is studied to aid expert specialization.
@@
  • Extensive experiments on UPMC Food-101, SNLI-VE and MM-IMDB datasets demonstrate state-of-the-art performance and parameter-efficiency for multimodal fusion.
2 Related Works
@@
Prompt Tuning for Transfer Learning. Prompt tuning involves learning continuous token embedding as additional input to a frozen pretrained model for transfer learning. This technique was first popularized in NLP [20, 19], then quickly introduced to CV society [11, 2, 25, 23] and multimodal learning [12, 14, 49, 47]. For either modality, researchers empirically report good transfer learning performance in low-shot adaptation, yet its performance is less comparable to fine-tuning when abundant training instances are available [20, 7]. Moreover, increasing prompt length quickly reaches performance saturation, and over-length prompts might have an negative impact on performance [44, 11, 16]. Recent theoretical analysis [32, 42] reveal that the expressiveness of vanilla prompt tuning is lower than fine-tuning and LoRA [10]. In this work, we scale up the expressiveness of vanilla prompting by employing a Mixture of Experts (MoE)-like design.
@@
Multimodal Fusion via Prompt. Apart from adapting models, prompts can also be used to fuse separately pretrained unimodal models for multimodal tasks. This is a different task from the commonly known “multimodal prompt tuning”, which usually refers to using prompts for transfer learning of multimodal pretrained foundation models. Frozen [39] first introduces a method where the visual representation is mapped as a few input tokens to query frozen language models (LM). PromptFuse and BlindPrompt [22] improved upon this by introducing tunable prompts to the LM for cross-model alignment. PICa [45] translate images into discrete text captions to prompt frozen LMs. These methods treat tokens from different modalities independently and lack cross-modal interaction. Recently, PMF [21] introduced an interactive prompt-tuning method, but their interactivity is built on the strong assumption that both encoders are white-box and in the same Transformer architecture. In this work, we further investigate how to efficiently utilize cross-modal features for prompting while maintaining high modularity by treating one modality as a black box.
@@
MoE in Transformers. MoE is a powerful technique to scale up models, including Transformers [18, 36, 31, 6]. The fundamental approach involves inserting MoE layers, composed of multiple feed-forward networks (FFNs) acting as experts, into the standard Transformer architecture. A router is learned to route each token embedding to the most suitable expert(s). Due to the significant computational cost associated with FFN forwarding, a sparse gating function is typically employed to limit the number of experts used per token [33, 36, 38]. Thanks to those designs, the MoE layer can efficiently scale up model capacity while maintaining computational efficiency. Inspired by previous FFN-based MoE, we explore scaling up the vanilla prompting with an MoE design.
3 Method
Our objective is to adaptively and efficiently fuse unimodal pretrained models for solving multimodal tasks via prompt tuning. In this section, we first review the working mechanism and the limitation of vanilla prompt tuning in Sec. 3.1. In Sec. 3.2, we explain how we achieve conditional prompt-tuning via a sequential pipeline and disentangled prompt tuning. In Sec. 3.3 we introduce the proposed MoPE, consisting of multiple prompt experts and a mutimodal router for expressiveness scaling. Finally, we introduce the regularization methods in Sec. 3.4.
@@ 3 Method
Figure 2: Architecture overview. We illustrate the proposed MoPE-based multimodal fusion method when applied to one frozen Transformer layer. a) We sequentially fuse two modalities, using modality y to guide the prompting of modality x through a three-type prompt disentanglement process. (b) MoPE is introduced, utilizing the multimodal pairing as a prior to instance-wisely generate the dynamic prompt; (c) Inside the multimodal router, we project the representation from each modalities. The concatenated feature (along channel dimension) is used to match the routing embedding paired with each experts for routing score calculation. Better viewed with color.
3.1 Preliminary: Vanilla Prompt Tuning
@@
Consider a Transformer [40] or its variants used to extract features from an embedded input sequence $\mathbf{x}^{0}\in\mathbb{R}_{s\times d_{x}}$, where $s$ is the sequence length and $d_{x}$ is the embedding dimension of the Transformer. Prompt-tuning [^1] In this paper we assume that soft prompts are applied to all layers, which corresponds to “VPT-deep” in CV and “prefix-tuning” in NLP. freezes all pre-trained model weights and optimize a small amount of continuous embeddings (i.e., prompts) $\mathbf{P}\in\mathbb{R}_{l\times d}$ inserted into the input, where $l$ is the number of prompts or “prompt length”. The input of the $i$-th layer layer $L^{i}$ could be denoted as:
@@
Succinctly, prompt tuning works by biasing the pretrained attention pattern in each Transformer layers [32]. As opposed to fine-tuning, this biasing has strictly limited expressiveness [32, 42], meaning that there are tasks that are un-learnable even with $l\to\infty$. The empirical expressiveness of vanilla prompting is even lower than its theoretical upper bound, due to competing optimization when $l>1$, as postulated by Petrov et al. [32]. In other words, although longer prompts provide the model with more capability, finding such a long prompt becomes challenging, and a suboptimal long prompt can lead to performance degradation. In this paper, we do not aim to increase the theoretical upper bound of prompt-tuning; instead, we approach the problem by narrowing the gap between its empirical and theoretical expressiveness.
3.2 Conditional Prompt Tuning for Multimodal Fusion
Previous prompt-based multimodal fusion methods [22, 21] optimize a global prompt that is shared across all instances. However, this approach neglects the interplay between different modalities. With paired multimodal input as a prior, we aim to condition the prompting of one modality on the other for more adaptive and efficient transfer learning.
Sequential Pipeline. A sequential pipeline for multimodal fusion is adopted to achieve prompt conditioning. Specifically, let $(x,y)$ represent a pair of multimodal inputs, and $\mathcal{E}_{y}$ be the encoder of modality $y$. Depending on the intrinsics of the specific task, a fusion direction could be assigned from the complementary modality to the main modality, and the proposed method will be used for tuning the main modality only. Without loss of generality, we consider $x$ to be the main modality in the following discussion.
@@ 3.2 Conditional Prompt Tuning for Multimodal Fusion
Prompt Disentanglement. Based on the sequential fusing pipeline, we disentangle the vanilla prompt vector $\mathbf{P}$ used in Eqn. 1 into three types of specialized prompts $[\mathbf{P}_{s},\mathbf{P}_{d},P_{m}]$. The static prompt $\mathbf{P}_{s}\in\mathbb{R}_{l\times d_{x}}$ is a globally-shared prompt vector that is non-conditional to input. To synthesize the dynamic prompt $\mathbf{P}_{d}\in\mathbb{R}_{l\times d_{x}}$, we first extract the global-level feature from the complementary modality $\mathcal{E}_{y}(y)=\psi_{y}\in\mathbb{R}_{d_{y}}$, and utilize a MoPE module $R(\cdot)$ to synthesize the instance-wise dynamic prompt (entailed in Sec 3.3). Additionally, we apply a lightweight mapper $f_{m}(\cdot)$ to map the complementary feature $\psi_{y}$ into a single prompt $P_{m}\in\mathbb{R}_{d_{x}}$, which injects fine-grained cross-modal information. In summary, the input of layer $L^{i}$ of $\mathcal{E}_{x}$ becomes:
$\mathbf{\hat{x}}^{i}=[x^{i-1}_{0},\underbrace{\mathbf{P}_{s},R(\psi_{y}),f_{m}(\psi_{y})}_{\text{Disentangled prompts}},\mathbf{T}^{i-1}]$ (2)
It is important to note that our three types of prompting serve as a plug-in module to replace the vanilla prompt vector of the main modality, while treating the encoder of the complementary modality $\mathcal{E}_{y}$ as a black-box for modularity. The whole process is illustrated in Fig. 2-(a).
@@
3.3 Mixture of Prompt Experts
As length-scaling has been proven ineffective due to quick saturation and over-length deterioration, we propose to fix the prompt length used in attention to suppress the negative impact of longer prompts. To inject more learnable parameters into the framework and increase the expressiveness of the prompts, we condition the prompts on multiple prompt experts for end-to-end training.
Prompt as Expert. We randomly initialize $k$ prompt experts and their associated routing embedding $\{(\mathbf{E}_{j},\mathbf{v}_{j})\}_{j=1}^{k}$ to be optimized end-to-end in each prompt-tuned Transformer layer $L^{i}$. For each pair, $\mathbf{E}_{i}\in\mathbb{R}_{l\times d}$ is the expert and $\mathbf{v}_{i}\in\mathbb{R}_{d_{r}}$ is the routing embedding, where $d_{r}$ is dimension of routing embedding.
Multimodal Expert Routing. In multimodal tasks, missing information from either modality can lead to ambiguous and incorrect results. Therefore, we route each instance based on representations from both modalities. Specifically, for layer $L^{i}$, we apply a layer-wise learned linear transformation $\mathbf{W}_{y}^{i}\in\mathbb{R}_{d_{y}\times d_{c}}$ to get the cross-modal routing embedding based on cross-modal feature $\psi_{y}$, where $d_{c}$ is the embedding dimension for cross-modal routing. We also utilize the global-level feature from previous layer $L^{i-1}$ to compute a inter-modal routing embedding through $\mathbf{W}_{x}^{i}\in\mathbb{R}_{d_{x}\times d_{i}}$, where $d_{i}$ is the inter-modal routing embedding, and $d_{i}+d_{c}=d_{r}$. Finally, we concatenate both embeddings along the embedding dimension to get an adjoint embedding $\mathbf{k}\in\mathbb{R}_{d_{r}}$, and dot product them with the routing embedding of all available experts. The routing score $\mathbf{r}$ is calculated by:
$\mathbf{k}=[\psi_{y}\mathbf{W}_{y}^{i},x_{0}^{i}\mathbf{W}_{x}^{i}],\quad\mathbf{r}_{j}=\frac{\operatorname{exp}(\mathbf{k}^{\top}\mathbf{v}_{j}/\tau+\epsilon)}{\sum_{n=1}^{k}\operatorname{exp}(\mathbf{k}^{\top}\mathbf{v}_{n}/\tau+\epsilon)}$ (3)
where $\tau=0.1$ is the temperature hyper-parameter, and $\epsilon$ is sampled noise [36]. The dynamic prompt is calculated by a convex combination of all experts at this layer according to the routing score:
@@ 3.3 Mixture of Prompt Experts
3.4 Regularizing Expert Routing
The proposed MoPE scales up vanilla prompt tuning by dividing the problem space into subspaces governed by specialized experts. However, we empirically found that a few experts tend to dominate across all instances during training, resulting in non-specialization. This phenomenon is self-reinforcing and is due to degenerated routing, where experts receiving high scores during initialization are exposed to larger gradient scales, causing them to be selected more frequently during optimization [5, 38]. In this section, we introduce methods to circumvent this issue and aid expert specialization.
Orthogonal Routing Embedding. Inspired by the analysis of degenerated routing, we freeze the routing embeddings to avoid self-reinforcement. We find that an orthogonal initialized [37] and non-learnable routing embedding $\{\mathbf{v}\}_{j=1}^{k}$ performs as well as (and even slightly better than) learned embeddings, while requiring fewer trainable parameters.
@@ 3.4 Regularizing Expert Routing
Importance Loss. To aid expert specialization, we add an additional importance loss [38, 36] to penalize dominant experts. For a batch of input $\mathbf{B}=\{(x_{0},y_{0}),(x_{1},y_{1}),...,(x_{b},y_{b})\}$, the importance of expert-$j$ is defined as the summed routing score in this batch. The importance loss is calculated as the mean coefficient of variation of all experts’ importance averaged across all layers::
@@
$\operatorname{Imp}(\mathbf{E}_{j})=\sum_{(x,y)\in\mathbf{B}}\mathbf{r}_{j},\quad\mathcal{L}_{imp}=\operatorname{stopgrad}\left(\left(\frac{\operatorname{std}(\{\operatorname{Imp}(\mathbf{E}_{j})\}_{j}^{k})}{\operatorname{mean}(\{\operatorname{Imp}(\mathbf{E}_{j})\}_{j}^{k})}\right)^{2};\gamma\right)$ (5)

### v2

@@ Abstract
Despite the demonstrated parameter efficiency of prompt-based multimodal fusion methods, their limited adaptivity and expressiveness often result in suboptimal performance compared to other tuning approaches. In this paper, we address these limitations by decomposing the vanilla prompts to adaptively capture instance-level features. Building upon this decomposition, we introduce the mixture of prompt experts (MoPE) technique to enhance the expressiveness of prompt tuning. MoPE leverages multimodal pairing priors to route the most effective prompt on a per-instance basis. Compared to vanilla prompting, our MoPE-based fusion method exhibits greater expressiveness, scaling more effectively with the training data and the overall number of trainable parameters. We also investigate regularization terms for expert routing, which lead to emergent expert specialization during training, paving the way for interpretable soft prompting. Extensive experiments across six multimodal datasets spanning four modalities demonstrate that our method achieves state-of-the-art results for prompt fusion, matching or even surpassing the performance of fine-tuning while requiring only 0.8% of the trainable parameters.
Introduction
Unimodal pre-trained models like BERT (Radford et al. 2019) and Swin Transformer (Liu et al. 2021b) have excelled in transferring capabilities to a wide range of tasks. By comparison, extending the pretraining-finetuning paradigm to multimodal applications involves additional complexities, particularly regarding flexibility and cost. First, the modality-paired pretraining can be inflexible due to its dependence on specific configurations. This can limit flexibility when there is a scarcity of certain modality pairs, or when new architectures emerge, necessitating resource-intensive retraining. In contrast, unimodal pretrained models are more readily available and offer greater flexibility, as they can be independently updated and combined for multimodal tasks. Second, as these foundation models scale in size, fine-tuning them becomes increasingly expensive, which further restricts their application to downstream tasks. To democratize foundation models, a compelling question arises: How can we efficiently combine separately pretrained unimodal models for multimodal tasks?
@@ Introduction
Figure 1: High-level motivation of MoPE-based multimodal fusion. (a) Vanilla prompt tuning learns a globally shared long prompt for all instances, which may not be optimal for each instance. (b) MoPE decomposes the global prompt into multiple short and specialized prompt experts to improve its adaptivity and expressiveness.
@@
While originally proposed for transfer learning (Li and Liang 2021; Jia et al. 2022; Zhou et al. 2022b), recent research has revealed that prompt tuning could also be a parameter-efficient paradigm for multimodal fusion (i.e., prompt fusion). Typically, this is achieved by representing one modality with several prompts, which are fed into the frozen Transformer of another modality to achieve fusion (Tsimpoukelli et al. 2021; Liang, Zhao, and Schütze 2022; Li et al. 2023; Yang et al. 2022b). However, directly using prompts for fusion could yield suboptimal results, particularly due to their limited adaptivity and scalability. Notably, while prompt tuning generally performs well in low-data regimes, it can be less effective when applied to full-shot training on the entire dataset with a challenging objective (Gao et al. 2022; Liang, Zhao, and Schütze 2022; Tsimpoukelli et al. 2021; Yang et al. 2022a; Lester, Al-Rfou, and Constant 2021). This reduced efficacy could be attributed to at least two factors. First, the vanilla prompt was originally developed for unimodal learning, which optimizes globally-shared prompts for all instances (Jia et al. 2022; Li and Liang 2021; Tsimpoukelli et al. 2021). When directly applied for prompt fusion (Tsimpoukelli et al. 2021; Li et al. 2023; Liang, Zhao, and Schütze 2022), it fails to adaptively exploit the multimodal interplay and can lead to sub-optimal representations for each input instance. Second, the relatively limited representation power (compared to fine-tuning) can lead to underfitting in multimodal datasets with a long-tail distribution and complex cross-modal mapping (Liu et al. 2021a; Petrov, Torr, and Bibi 2023; Wang et al. 2023).
@@
To address these challenges, increasing the number of learnable prompts, known as “length-scaling”, appears to be a straightforward solution. However, numerous studies have shown that the performance gains from length-scaling quickly reach saturation in both transfer-learning (Gao et al. 2022; Lester, Al-Rfou, and Constant 2021) and fusion settings (Liang, Zhao, and Schütze 2022; Tsimpoukelli et al. 2021; Yang et al. 2022a). Furthermore, over-length prompts may even lead to worse results (Jia et al. 2022; Yang et al. 2022a; Khattak et al. 2023; Li and Liang 2021; Kim et al. 2023). Recent theoretical analyses have substantiated these empirical observations, highlighting the limited expressiveness of prompt tuning (Wang et al. 2023; Petrov, Torr, and Bibi 2023). Overall, evidence from both empirical and theoretical analyses reveals that simply increasing prompt length is insufficient to address the limitations of prompt fusion.
To improve the adaptivity and expressiveness of prompt fusion, we propose decomposing the unified, long prompt into multiple short prompts with specialized responsibilities, as depicted in Fig. 1. We introduce an instance-wise adaptive prompt decomposition that separates the vanilla prompt into static, dynamic, and mapped prompts to adaptively capture global and local features. The foundation of our dynamic prompt generation lies in the proposed MoPE (Mixture of Prompt Experts) design, where we maintain a fixed prompt length in the self-attention and enhance expressiveness by learning multiple prompt experts and a multimodal router to select the most suitable prompt for each instance.
We systematically compare our MoPE-based fusion with previous prompt fusion methods on a total of six datasets, and demonstrate consistently superior performance with higher parameter-efficiency. Furthermore, we showcase the superior scalability of our method both in terms of parameter and data. Our analysis indicates that increasing the number of experts (i.e., “expert-scaling”) is more scalable than length-scaling, with monotonic performance gains and the avoidance of performance deterioration with an overly-long prompt. Our method also scales better with more training data. By introducing regularization terms for expert routing, we observe the emergence of specialized prompt experts after end-to-end training, resulting in high interpretability. Our key contributions are summarized as follows:
  • We propose an instance-wise adaptive prompt decomposition, which decomposes the vanilla prompt into static, dynamic and mapped prompt.
@@
  • A combination of regularization terms is studied to aid specialization of prompt experts.
@@
  • Extensive experiments on six datasets spanning four modalities demonstrate state-of-the-art performance and parameter efficiency for prompt-based multimodal fusion.
Related Works
@@
Prompt Tuning for Transfer Learning. Prompt tuning (Li and Liang 2021; Lester, Al-Rfou, and Constant 2021) learns continuous token embeddings as additional input to a frozen pretrained model for transfer learning of Transformer-based models. It is widely used in various modalities (Jia et al. 2022; Bahng et al. 2022; Liu et al. 2022; Lin et al. 2023; Jiang, Liu, and Chen 2023; Khattak et al. 2023; Zhou et al. 2022a). A common observation is good transfer learning performance in low-shot scenarios, yet its performance is less comparable to fine-tuning when abundant training instances are available (Li and Liang 2021; Gao et al. 2022). Moreover, increasing prompt length quickly reaches performance saturation, and over-length prompts might lead to worse results (Yang et al. 2022a; Jia et al. 2022; Kim et al. 2023). Recent theoretical analyses (Petrov, Torr, and Bibi 2023; Wang et al. 2023) reveal that the expressiveness of prompt tuning is lower than that of fine-tuning. In this work, we tackle this challenge by scaling up the expressiveness of prompts with a Mixture of Experts (MoE)-like design.
@@
Multimodal Fusion via Prompt. Apart from model adaptation, prompts can also be used to fuse separately pretrained unimodal models for multimodal tasks. Frozen (Tsimpoukelli et al. 2021) first introduced a method where the visual representation is mapped as a few input tokens to query frozen language models (LMs). PromptFuse and BlindPrompt (Liang, Zhao, and Schütze 2022) improved upon this by introducing tunable prompts to the LM for cross-modal alignment. PICa (Yang et al. 2022b) translates images into discrete text captions to prompt frozen LMs. These methods treat tokens from different modalities independently and lack cross-modal interaction. Recently, PMF (Li et al. 2023) introduced an interactive prompt fusion method for the vision and text modalities, building on the strong assumption that both encoders are white-box and have the same Transformer architecture. In this paper, we investigate how to adaptively interact with two or more modalities with minimal assumptions about the model architecture.
@@
MoE in Transformers. MoE is a computationally efficient technique to scale up models, including Transformers (Lepikhin et al. 2020; Riquelme et al. 2021; Mustafa et al. 2022; Fedus, Zoph, and Shazeer 2022). The fundamental approach involves inserting MoE layers, usually composed of multiple feed-forward networks (FFNs) acting as experts, into the standard Transformer architecture. A router is learned to route each token embedding to the most suitable expert(s) for reducing computational overhead (Qin and Eisner 2021; Riquelme et al. 2021; Shazeer et al. 2017). In this paper, we do not focus on reducing computational cost; instead, we introduce the MoE design into prompt fusion to scale up its adaptiveness and expressiveness.
Method
@@
Figure 2: Architecture overview. a) A sequential fusion pipeline is used, using modality $\mathbb{Y}$ to guide the prompting of modality $\mathbb{X}$ through our prompt decomposition. (b) We introduce MoPE, which utilize the multimodal pairing as a prior to route the most suitable dynamic prompt for each instance; (c) Inside the multimodal router, we project the representation from each modalities. The concatenated embedding is used to match the routing embedding paired with each experts for routing score calculation. Better viewed with color.
Preliminary: Vanilla Prompt Tuning
@@
Consider a Transformer (Vaswani et al. 2017) or its variants used to extract features from an embedded input sequence $\mathbf{x}^{0}\in\mathbb{R}_{s\times d_{x}}$, where $s$ is the sequence length and $d_{x}$ is the embedding dimension. Prompt-tuning freezes all pre-trained parameters and optimize a small amount of continuous embeddings (i.e., prompts) $\mathbf{P}\in\mathbb{R}_{l\times d}$ concatenated to the input of each layer, where $l$ is length of the prompt. The input of the $i$-th layer layer $L^{i}$ could be denoted as:
@@
Succinctly, prompt tuning works by biasing the pretrained attention pattern in each Transformer layers (Petrov, Torr, and Bibi 2023). As opposed to fine-tuning, this biasing has strictly limited expressiveness in theory (Petrov, Torr, and Bibi 2023; Wang et al. 2023), meaning that there are tasks that are un-learnable even with $l\to\infty$. The empirical expressiveness of vanilla prompting is even lower than its theoretical upper bound, due to competing optimization when $l>1$ (Petrov, Torr, and Bibi 2023). In other words, although longer prompt gives more trainable parameters, finding such a long prompt becomes challenging, and a suboptimal long prompt can lead to performance degradation instead. In this paper, we do not aim to increase the theoretical upper bound of prompt-tuning; instead, we approach the problem by narrowing the gap between its empirical and theoretical expressiveness.
Instance-wise Adaptive Prompt Decomposition
Previous prompt fusion methods (Liang, Zhao, and Schütze 2022; Yang et al. 2022b) optimize a global prompt that is shared across all instances, neglecting the interplay between different modalities and instance-specific features. With paired multimodal input as a prior, we aim to instance-wisely condition the prompting of one modality on the other(s) for better adaptivity. To achieve this goal, we first adopt a sequential fusion pipeline. Let $x\in\mathbb{X},y\in\mathbb{Y}$ be a pair of multimodal inputs, and $\mathcal{E}_{\mathbb{X}}$, $\mathcal{E}_{\mathbb{Y}}$ be the encoder of each modality. Depending on the task intrinsic, we assign a fusion direction where the encoding of main modality $\mathbb{X}$ is conditioned on representation of complementary modality(ies) $\mathbb{Y}$.
@@
Building upon the sequential pipeline, we decompose the vanilla prompt vector $\mathbf{P}$ used in $\mathcal{E}_{\mathbb{X}}$ into three types of specialized prompts $[\mathbf{P}_{s},\mathbf{P}_{d},P_{m}]$. The static prompt $\mathbf{P}_{s}\in\mathbb{R}_{l\times d_{x}}$ is a globally-shared prompt vector that is shared for all instances. The dynamic prompt $\mathbf{P}_{d}\in\mathbb{R}_{l\times d_{x}}$ are adaptive to different instances. To synthesize it, we first encode global-level feature from the complementary modality $\psi_{y}=\mathcal{E}_{\mathbb{Y}}(y)\in\mathbb{R}_{d_{y}}$, and utilize a MoPE module $R(\cdot,\cdot)$ to generate the dynamic prompt. Additionally, we apply a lightweight mapper $f_{m}(\cdot)$ to map the complementary feature into a single prompt $P_{m}\in\mathbb{R}_{d_{x}}$, which injects fine-grained cross-modal information. In summary, the input of layer $L^{i}$ of $\mathcal{E}_{\mathbb{X}}$ becomes:
$\mathbf{\hat{x}}^{i}=[x^{i-1}_{0},\underbrace{\mathbf{P}_{s},R(x_{0}^{i-1},\psi_{y}),f_{m}(\psi_{y})}_{\text{Decomposed prompts}},\mathbf{T}^{i-1}]$ (2)
The whole process is illustrated in Fig. 2-(a).
@@
Mixture of Prompt Experts
To mitigate the negative impact on optimizing a long prompt and to improve the adaptiveness of dynamic prompt, we propose to learn multiple short prompts with a MoE-like design. To be more specific, we use MoPE module in each prompt-tuned Transformer layers $L^{i}$ to generate the dynamic prompt in a scalable way. A MoPE module consists of a router, $k$ prompt experts and their associated routing embeddings $\{(\mathbf{E}_{j},\mathbf{k}_{j})\}_{j=1}^{k}$ where $\mathbf{E}_{i}\in\mathbb{R}_{l\times d}$ is an expert, $\mathbf{k}_{i}\in\mathbb{R}_{d_{r}}$ is its routing embedding, and $d_{r}$ is the dimension of routing embedding.
We route each instance based on representations of all modalities, as illustrated in Fig. 2-(b,c). The router is parameterized as two layer-specific linear transformations $\mathbf{W}_{y}^{i}\in\mathbb{R}_{d_{y}\times d_{c}}$ and $\mathbf{W}_{x}^{i}\in\mathbb{R}_{d_{x}\times d_{i}}$, where $d_{c}$, $d_{i}$ are the dimension of cross-modal and inter-modal routing embedding, respectively, and $d_{i}+d_{c}=d_{r}$. The cross-model embedding is projected from $\psi_{y}$, while inter-model embedding is calculated over the global-level feature (typically represented as the [CLS] token) from previous layer $L^{i-1}$. Finally, we concatenate both embeddings to get an multimodal query embedding $\mathbf{q}\in\mathbb{R}_{d_{r}}$, and compute the dot product with the routing embeddings of all available experts. The routing score $\mathbf{r}$ is calculated by:
$\mathbf{q}=[\psi_{y}\mathbf{W}_{y}^{i},x_{0}^{i-1}\mathbf{W}_{x}^{i}],\quad\mathbf{r}_{j}=\frac{\operatorname{exp}(\mathbf{q}^{\top}\mathbf{k}_{j}/\tau+\epsilon)}{\sum_{n=1}^{k}\operatorname{exp}(\mathbf{q}^{\top}\mathbf{k}_{n}/\tau+\epsilon)}$ (3)
where $\tau=0.1$ is the temperature hyper-parameter, and $\epsilon$ is sampled noise (Riquelme et al. 2021). When there are more than one complementary modality, the additional embedding could be easily extended by learning additional projections. The dynamic prompt is obtained by a convex combination of all experts at this layer according to the routing score:
@@
Regularizing Expert Routing
We empirically find that a few experts tend to dominate across all instances during training, a phenomenon that has also been observed in previous MoE methods (Eigen, Ranzato, and Sutskever 2013; Shazeer et al. 2017). In this section, we introduce methods to circumvent degenerated routing and aid expert specialization.
Orthogonal Routing Embedding. To avoid the expert being selected in initialization being exposed to a larger gradient signal, we freeze the routing embeddings $\{\mathbf{k}\}_{j=1}^{k}$ to suppress self-reinforcing expert domination. Our finding indicate that an orthogonal initialized (Saxe, McClelland, and Ganguli 2013) and non-learnable routing embedding outperform learned one, while requiring fewer trainable parameters.
@@
Importance Loss. To further aid expert specialization, we add an additional importance loss (Shazeer et al. 2017; Riquelme et al. 2021) to penalize dominant experts. For a batch of input $\mathbf{B}=\{(x_{0},y_{0}),(x_{1},y_{1}),...,(x_{b},y_{b})\}$, the importance of expert-$j$ is defined as the summed routing score in this batch, $\operatorname{Imp}(\mathbf{E}_{j})=\sum_{(x,y)\in\mathbf{B}}\mathbf{r}_{j}$. The importance loss is calculated as the mean coefficient of variation of all experts’ importance averaged across all layers:
@@
$\mathcal{L}_{imp}$ $=\sum_{All\ layers}\operatorname{\sigma}\left(\left(\frac{\operatorname{std}(\{\operatorname{Imp}(\mathbf{E}_{j})\}_{j}^{k})}{\operatorname{mean}(\{\operatorname{Imp}(\mathbf{E}_{j})\}_{j}^{k})}\right)^{2};\gamma\right)$ (5)